OpenStudy (testing123guest):

Multiply and answer in scientific notation: (3.8 x 10^8)(6.9 x 10^5)

OpenStudy (testing123guest):

@reneesweet17

OpenStudy (anonymous):

ok first what is \[10^8\] solve that first it just like the other one just more steps

OpenStudy (testing123guest):

10⁸ = 10×10×10×10×10×10×10×10 = 100000000

OpenStudy (testing123guest):

bro u left me

OpenStudy (anonymous):

nice ok so next right that down on paper and do the same with \[10^5\]

OpenStudy (testing123guest):

10⁵ = 10×10×10×10×10 = 100000

OpenStudy (anonymous):

ok good now change \[3.8\] into a regaler number

OpenStudy (anonymous):

did u get it

OpenStudy (anonymous):

@Testing123guest

OpenStudy (testing123guest):

ya it was 2.622 * 10^14

OpenStudy (testing123guest):

@reneesweet17

OpenStudy (anonymous):

ok so u have the where did u get the 10^14

OpenStudy (anonymous):

where did u get the 10^14